From GlueXWiki
Jump to: navigation, search

January 29, 2015 FDC+CDC meeting


  1. Instructions for Bluejeans meeting connection
  2. FDC meeting ID: 290664653
  3. To join via a Web Browser, go to the page [1] https://bluejeans.com/290664653.


  1. Collaboration meeting agenda
  2. Gas puzzle
  3. CDC update (Mike)
  4. FDC update
  5. Electronics (Fernando, Cody)
    • CDC status (Nick, Chris)
  6. Engineering (Dave)
  7. Other


Participants: Curtis, Naomi (CMU), Eugene, Mike, Beni, Dave, Chris, Nick, Simon, Cody, Vlad, Sergey, Fernando, and Lubomir (JLab).

Collaboration meeting agenda

- the agenda seams to be finalized (see link above).

Gas puzzle

- Vlad and Lubomir prepared a small setup for the CO2 measurements (pictures). The CO2 sensor was calibrated in the EEL126 with a certified gas mixture of Ar/CO2 with 59.66% CO2. Then we moved the setup to the hall and connected it at the same place where the oxygen sensor is supposed to be - right before the exhaust bubblers - where, using solenoid valves you can choose different chambers. We did short measurements with the CDC and the first FDC package - the results (including calibration) are summarized in the linked ppt file: CDC has 41.4%, FDC has 55.3% CO2 +/-0.3% RMS. The CDC measurements showed some variations but monitoring over longer period of time is needed.

- Right after these measurements, the software could no longer recognize the sensor and we were not able to do more measurements. Lubomir is contacting now the company to resolve the problem. We had also another issue with the company: the plastic adapter cup on the top of the sensor had holes drilled in a wrong place (picture attached), but Scot fixed it.

- Noami did measurements with the prototype at CMU for gas mixtures with 50,55 and 60% Ar, results for the most probable amplitudes and maximum drift times are plotted by Beni. Comparing to the results at JLab one estimates the gas in the Hall to be ~57.5% Ar, very similar when using amplitudes or drift times and also close to the CO2 measurements: 100% - 41.4%CO2 -0.33% propanol = 58.3% Ar. Keep in mind there are some differences between the prototype and the CDC in the wire diameter (10%) and in the straw diameter. Also the effect of the alcohol at JLab is not taken into account.

- In summary, even with some uncertainty, we have in the CDC ~42% CO2 instead of 50%, and in the FDC the CO2 is 55% instead of 60%. It is not clear at all why this is happening. Both, CDC and FDC results can be explained by 20-22% increase of the argon. Beni thinks it might be due to the AR/CO2 separation (due to the density difference) in the mixing tank and proposes to use a heater to create convection. The plan will be to test all this with the CDC including the effect of the alcohol.

- Lubomir used the measured CO2 values for Garfield simulations of the time-to-distance functions, assuming 0.33% propanol (from vapor pressure estimates) for both FDC and CDC (plots attached). Comparing with the measurement for the CDC the gas calculated by Garfield is a little faster, i.e. needs some more CO2. For the FDC, Garfield describes very well the measurements, except in the region of 0.5-1.5 mm which can't be fixed by changing only the AR/CO2 ratio, i.e. the alcohol may play a role there. As Naomi found, Garfield has data only for isopropanol, which is used for both propanol and isopropanol.

FDC update

- Lubomir showed results for strip resolutions after aligning one package: estimated from tracking it is ~250 microns, while from reconstructing the wire position it should be ~170 microns. This may mean that better alignment of all the cells is needed. At the same time it may be related to some structures seen in the last plot above: the difference between the cathode reconstructed and the nominal wire position as a function of x and y averaged over all the hits for each 2d-point. The dips and peaks on the plot rage by up to 200 microns and this structure is very similar for all the cells in the package. It's impossible to have such deviations in the wire positions, so it must be coming from the strip positions; needs further investigation.

- Wire resolution: estimated from tracking the over-all resolution is ~200 microns. Plotted as a function of the distance to the wire it is somewhat worse than the results with cosmics at EEL126 for small distances. This may indicate worse timing resolution coming from t0 uncertainties when using the start counter.


- Fernando: CDC is ready for testing, all HVBs have been installed and the HV cables re-routed. However, the noise is still there. CAEN finished the prototype of the chassis with the HV filter. It will be tested here and then decide about the production.

- Cody installed the latest f125 firmware and started tests: it works as expected, except in block mode when the triggers are too close (Beni). Issues with some of the channels turned out to be software problem: bad DAC values.

- One of the signal cables on the CDC turned out to be shorter than the rest, that's why the signals from this card were coming earlier.


- Dave: the thremocouple for the alcohol bubbler is now working and the temperature is sent as an EPICS variable. The power switch-off box for the chiller will be tested with a motor. The chiller has been received by the company and soon they will have an estimate about the time needed for the repair.